Nina is four times as old as Sam. Next year, she will be only three times as old as Sam. How old are the kids?

Found 2 solutions by math_helper, ikleyn:
Answer by math_helper(2461)

The question is missing information.
Also, If NinasAge = 3*SamsAge then NinasAge+1 can not be SamsAge*4, unless we are talking about fractions of one year.
Edit: Doh! I should've worked it out on paper a bit, then I would've saw that it is possible, 'kids' being the keyword.

Answer by ikleyn(52803)

.
N = 4*S,          (1)
N+1 = 3*(S+1).    (2)


===========>  


4*S + 1 = 3*S + 3  ====>  S = 2.


Answer.  Sam is 2 years old. Nina is 4*2 = 8 years old.


Everything is OK with the input data.

---------------
There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.